OpenStudy (madgirlwithabluebox):

please help? solve log(2x+3)=3

OpenStudy (madgirlwithabluebox):

A. 500 B. 997 C. 1/2 D. 997/2

OpenStudy (madgirlwithabluebox):

i will medal if someone explain this to me :(

OpenStudy (anonymous):

\[\log \left( 2x+3 \right)=3\] \[2x+3=10^3=1000\] 2x=1000-3=997 x=997/2

OpenStudy (madgirlwithabluebox):

can you help with one more because i still dont understand how to do this..